DIRC Operations Weekly Report

Period: Saturday 2006/07/08 - Friday 2006/07/14

Oncall: Rolf Andreassen, Jerome Bequilleux
Fast Monitoring: Bradley Wogsland, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060701.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060623 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 11.2*1033 (previous week: 10.2*1033)

BaBar integrated lumi: 3.98 fb-1 (previous week: 2.831 fb-1)

Run 5b PEP delivery: 62.49 fb-1

Run 5 total delivery: 124.75 fb-1 Run 5 total BaBar recorded: 120.24 fb-1 including 9.82 fb-1 off-peak data

PEP activities this week:

  • Saturday 24th/Sunday 25th : 2 major problems with PEP, water pump in region 12 + SK5 power supply -> PEPII ran ~50% time during the week-end. However, new luminosity record on Sunday (11.035*10^33)
  • Monday 26th : maintenance work on UPS system; problem with inverter drive boards remaining (another maintenance needed when spere boards will be available; hopefully it won' t disturb data taking). Began at 10am; beams back at 10:30pm.
  • Tuesday 27th : UPS spare available; maintenance work scheduled.
  • Wednesday 28th : new luminosity record at 11.24*10^33 (swing shift)
  • Thursday 29th : UPS work finished successfully; UPS is in normal operation again (intervention of Jerry to disable the SOB dump valve and prevent us from from any problem with SOB).

    PEP activities planned next week :

  • Foreseen work opportunities (PEP down periods):

  • BaBar activities/issues this week:

  • DIRC activities

    See DIRC logbook for details
  • June 24th: IR2 paged Rolf because bin 11 in FX plot was too high (In fact, just due due difficulties in injecting from PEP).
  • June 26th: UPS replacement. 2 problems occured:* DRC-MON did not reboot properly (solved by Ray); * GMB status for the gaz system showed a red zero (fixed by Matt).
  • June 29th: problem with HV crate 3, board #3, powering the 8 outer HV channels of sector 4. The decision was taken to replace the board. Questions about how to deal with crate 3.

    Downtime due to the DIRC: 17 minutes.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Manpower situation

    Links